OpenStudy (anonymous):

Write a proportion that could be used to solve for each variable. Then solve. 18 ounces in 6 bottlesx ounces in 24 bottles a. 18/6=x/24; x=72 b. 18/6=x/24; x=82 c. 6/18=x/24; x=8 d. 18/6=24/6;x=9

OpenStudy (anonymous):

@heather040200

OpenStudy (jh3power):

18/6=x/24 A x=18*4=72

OpenStudy (jh3power):

A

OpenStudy (anonymous):

thank you.